OpenStudy (anonymous):

What is the radius of a circle with a circumference of 23.55 meters? Use 3.14 for pi show work.

OpenStudy (campbell_st):

well the circumference is \[C = 2 \times \pi \times r\] so if you substitute your know information \[23.55 = 2 \times 3.14 \times r\] you can solve the equation for r

OpenStudy (anonymous):

Oh Thanks :D